OpenStudy (anonymous):

Okay, so if you don't like math then don't try to answer this. I've spent the last hour doing it. ∫3/(x^3−1)dx

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

\[\int\limits_{}^{}3/(x^3-1)dx\]

OpenStudy (anonymous):

is the '/' on purpose?

OpenStudy (anonymous):

Yeah, thats division.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I get lost when I'm at log|x-1|-\[\int\limits_{}^{}\frac{ x+2 }{ x^2+x+1}\]

OpenStudy (anonymous):

ln|x-1|*

OpenStudy (jtvatsim):

So you used partial fractions to get to this point correct?

OpenStudy (anonymous):

Yeah.

OpenStudy (jtvatsim):

Alright, so let me double check that on my end.

OpenStudy (anonymous):

I got

OpenStudy (jtvatsim):

Cool. I checked over here and it looks correct so far.

OpenStudy (jtvatsim):

Still thinking here... obviously, this is a tricky one. :)

OpenStudy (amistre64):

what part is losing you?

OpenStudy (amistre64):

\[\frac{ x+2 }{ x^2+x+1}\] \[\frac{ (x+1)+1 }{ x^2+x+1}\] what is the derivative of the bottom?

OpenStudy (anonymous):

Let u=x^2+x+1 and du=2x+1dx, \[\frac{ x+2 }{ x^2+x+1 }=\frac{ 1 }{ 2 }*\frac{ 2x+1 }{ x^2+x+1 }+\frac{ 3 }{ 2 }*\frac{ 1 }{ x^2+x+1 }\]

OpenStudy (anonymous):

I've gotten about up to that point, so you can say \[\int\limits_{}^{}\frac{ 1 }{ 2 }*\frac{ du }{ u }=\frac{ 1 }{ 2 }\ln|u|+C=\frac{ 1 }{ 2 }\ln|x^2+x+1|+C\]

OpenStudy (amistre64):

fine, we can do that too :) \[\frac12~\frac{ 2(x+1) }{ x^2+x+1}\] \[\frac12~\frac{ 2x+2) }{ x^2+x+1}\] \[\frac12~\frac{ (2x+1)+1 }{ x^2+x+1}\]

OpenStudy (anonymous):

But then \[\frac{ 3 }{ 2 } \int\limits_{}^{}\frac{ 1 }{ x^2+x+1 }dx\]

OpenStudy (amistre64):

completing the square on the bottom might be useful ... been awhile tho

OpenStudy (amistre64):

the other option is to just decompose it over the complex plane

OpenStudy (anonymous):

Oh okay, so its \[\frac{ 1 }{ (x+.5)^2+.75 }\] ? and now I'm pretty lost. It looks like maybe you can do some trig substitution but I can't really tell.

OpenStudy (amistre64):

what are your inverse trig derivatives? reviewing them might help out ... tan^-1 rings a bell

OpenStudy (anonymous):

Dont wanna type them all out so : http://tutorial.math.lamar.edu/Classes/CalcI/DiffInvTrigFcns_files/eq0044M.gif

OpenStudy (amistre64):

y = tan^-1 (x) tan(y) = x y' sec^2(y) = 1 y' = 1/sec^2(y) y' = 1/(tan^2(y)+1) y' = 1/(tan^2(tan^-1(x))+1) y' = 1/(x^2+1)

OpenStudy (amistre64):

typing them out helps to keep them in memory .... pauls site wont always be available :)

OpenStudy (anonymous):

True :P

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ (x+.5)^2+.75 }dx=\int\limits_{}^{}\frac{ \sqrt3/2*\sec^2(\theta) }{ 3/4*\tan^2\theta+1 }\]

OpenStudy (anonymous):

I think.... ._.

OpenStudy (amistre64):

im pretty sure im making a mess lol, but heres my thought process if we go this archaic route 1 -------------- (x+1/2)^2 + 3/4 4/3 ---------------- 4/3(x+1/2)^2 + 1 4/3 ---------------- (16/9 (x+1/2))^2 + 1 assuming we can work out some tan inverse y = K tan^-1(A(x+B)) y/K = tan^-1(A(x+B)) tan(y/K) = A(x+B) y' sec^2(y/K)/K = A y' sec^2(y/K) = KA y' (tan^2(y/K)+1) = KA y' ((A(x+B))^2+1) = KA y' = KA ------------ (A(x+B))^2+1 ------------------- A = 16/9, B=1/2 16K/9 = 4/3 16K = 12 K = 3/4

OpenStudy (anonymous):

Omg, \[\tan^2\theta+1=\sec^2\theta, so \it simplifies \to \int\limits_{}^{}\frac{ 2 }{ \sqrt3 }d \theta\]

OpenStudy (amistre64):

you are on the right track yes http://www.wolframalpha.com/input/?i=integrate+1%2F%28x^2%2Bx%2B1%29+dx

OpenStudy (amistre64):

mine went someplace off the wall lol

OpenStudy (amistre64):

4/3 sqrts to get inside the ^2 ...

OpenStudy (amistre64):

2/sqrt(3) not 16/9

OpenStudy (amistre64):

A = 2/sqrt(3) 2K/sqrt(3) = 4/3 2sqrt(3) K = 4 K = 2/sqrt(3) thats better on my end

OpenStudy (anonymous):

\[= \frac{ 2 }{ \sqrt3 } \theta+C, \theta=\arctan(\frac{ 2 }{ \sqrt3 }(x+.5)) \implies \frac{ 2 }{ \sqrt3 }\arctan(\frac{ 2 }{ \sqrt3}(x+.5))+C=\int\limits_{}^{}\frac{ 1 }{ x^2+x+1 }dx\]

OpenStudy (amistre64):

now me and the wolf agree: y = K tan^-1 (A(x+B)) \[y=\frac{2}{\sqrt3}~\tan^{-1}[\frac{2}{\sqrt3}(x+\frac12)]\]

OpenStudy (anonymous):

This answer is going to be really ugly in the end

OpenStudy (anonymous):

Yeah, I agree, then we just put it all together and its \[\log|x-1|-0.5\log|x^2+x+1|-\sqrt3\arctan(\frac{ 2 }{ \sqrt3 }(x+.5)+C\] I dont think that simplifies any further

OpenStudy (amistre64):

the uglier the better :)

OpenStudy (amistre64):

you got it

OpenStudy (anonymous):

Jeez, that was tough.

OpenStudy (amistre64):

nah .. it was "interesting" lol

OpenStudy (anonymous):

Lol

OpenStudy (amistre64):

good luck, its supper time

OpenStudy (anonymous):

Thank you btw.